# Fast convergence of Frank-Wolfe algorithms on polytopes

### v3

@@ Abstract
We provide a template to derive convergence rates for the following popular versions of the Frank-Wolfe algorithm on polytopes: vanilla Frank-Wolfe, Frank-Wolfe with away steps, Frank-Wolfe with blended pairwise steps, and Frank-Wolfe with in-face directions. Our template shows how the convergence rates follow from two affine-invariant properties of the problem, namely, error bound and extended curvature. These properties depend solely on the polytope and objective function but not on any affine-dependent object like norms. For each one of the above algorithms, we derive rates of convergence ranging from sublinear to linear depending on the degree of the error bound.
@@ 1.1 Outline
The expression ${\mathcal{O}}(t^{1/(2\theta-1)})$ in the above template is to be interpreted as linear rate of convergence in the special case $\theta=1/2$. The degree $\theta\in[0,1/2]$ of the error bound property yields a rate of convergence ranging from the sublinear rate ${\mathcal{O}}(t^{-1})$ when $\theta=0$ to a linear rate when $\theta=1/2$. The main constant hidden in the ${\mathcal{O}}(t^{1/(2\theta-1)})$ expression is the ratio $\mu^{2\theta}/L$ where $L$ and $\mu$ are parameters associated to the curvature and error bound properties. As we detail in Section 2, the error bound property is defined in terms of a distance function $\mathfrak{d}:\mathcal{C}\times\mathcal{C}\rightarrow[0,1]$. The suitable distance function for each version of the Frank-Wolfe algorithm is as summarized in Table 1. The radial $\mathfrak{r}$, vertex $\mathfrak{v}$, and face $\mathfrak{f}$ distance functions are defined in Section 3.
@@ 3 Algorithms
3.1 Vanilla Frank-Wolfe algorithm
@@
Functions of the form $\mathfrak{r}(\cdot,{\mathbf{x}})$ are somewhat intuitive and easy to visualize as they correspond to the gauge function of the set $\mathcal{C}-{\mathbf{x}}$. However, for our analysis functions of the form $\mathfrak{r}({\mathbf{y}},\cdot)$ are more relevant. Figure 4 displays visualizations of the function $\mathfrak{r}({\mathbf{y}},\cdot)$ on $\mathcal{C}$ for $\mathcal{C}=[0,1]\times[0,1]$ and various choices of ${\mathbf{y}}\in\mathcal{C}$. Figure 4(c) shows that when ${\mathbf{y}}$ lies on the relative boundary of $\mathcal{C}$, it holds that $\mathfrak{r}({\mathbf{y}},{\mathbf{x}})=1$ for every ${\mathbf{x}}\in\mathcal{C}$ with ${\mathbf{x}}\neq{\mathbf{y}}$. Figure 4(a) and Figure 4(b) show that a similar phenomenon occurs as ${\mathbf{y}}$ approaches the relative boundary of $\mathcal{C}$.
@@
For either of the line-search or short-step rules, the following progress bound holds for every iteration of Algorithm 1:
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle^{2}}{2L}.$ (3.1)
Our convergence analysis of FW relies on (3.1), and thus yields identical convergence rates for line-search and short-step.
@@
    $f({\mathbf{x}}^{(t)})-f^{*}\leq(f({\mathbf{x}}^{(0)})-f^{*})\left(1-\frac{\mu}{2L}\right)^{t}.$ (3.2)
  • When $\theta\in[0,1/2)$, for all $t\in\mathbb{N}$ the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(0)})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}t}{2L}\right)^{\frac{1}{2\theta-1}}.$ (3.3)
@@
Proof of Theorem 3.2. The $L$-curvature and the choice of step-size imply that (3.1) holds. Thus, Lemma 3.3 implies that for each $t\in\mathbb{N}$
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle^{2}}{2L}\leq f({\mathbf{x}}^{(t)})-\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{2(1-\theta)}}{2L}.$
Hence,
$f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\frac{\mu^{2\theta}}{2L}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}\right).$
Thus (3.2) readily follows when $\theta=1/2$. On the other hand, (3.3) follows by applying Lemma 2.6 with $p=1-2\theta$, $\sigma_{t}=\frac{\mu^{2\theta}}{2L}$, and $\beta_{t}=f({\mathbf{x}}^{(t)})-f^{*}$.
@@ 3.2 Away-step Frank-Wolfe algorithm and blended pairwise Frank-Wolfe algorithm
    For $t\in\mathbb{N}$ with $t_{0}\leq t\leq|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$ the following holds:
@@
    Finally, for $t\in\mathbb{N}$ with $t>|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$, the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(t_{0})})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}}{8L}\cdot\left\lceil\frac{t-|S({\boldsymbol{\lambda}}^{(t_{0})})|-t_{0}+1}{2}\right\rceil\right)^{\frac{1}{2\theta-1}}.$ (3.14)
@@
Proof of Theorem 3.6. The choice of ${\mathbf{d}}^{(t)}$ implies that for $t\in\mathbb{N}_{\geq 0}$
$\langle\nabla f({\mathbf{x}}^{(t)},{\mathbf{d}}^{(t)}\rangle\leq\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}^{(t)}-{\mathbf{a}}^{(t)}\rangle}{2}<0$ (3.17)
@@
    $f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{1}{2},\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}}{8L}\right\}\right).$ (3.19)
@@
$f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{1}{2},\frac{\mu}{8L}\right\}\right).$
@@
Thus  (3.13) holds for $t_{0}\leq t\leq|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$ and Lemma 2.6 implies (3.14) for $t>|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$.
@@ 3.3 In-face Frank-Wolfe algorithm
    And for $t\in\mathbb{N}_{\geq t_{0}}$, the following sublinear convergence rate holds:
@@
Proof of Theorem 3.12. This proof is similar that of Theorem 3.6 and thus we are deliberately succinct. The choice of ${\mathbf{d}}^{(t)}$ implies that for all $t\in\mathbb{N}_{\geq 0}$
$\langle\nabla f({\mathbf{x}}^{(t)},{\mathbf{d}}^{(t)}\rangle\leq\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}^{(t)}-{\mathbf{a}}^{(t)}\rangle}{2}<0$ (3.29)
@@
    $f({\mathbf{x}}^{(t+1)})-f^{*}\leq\left(1-\min\left\{\frac{1}{2},\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}}{8L}\right\}\right)(f({\mathbf{x}}^{(t)})-f^{*}).$
@@ 3.3.1 In-face FW for polytopes in standard form
    For $t\in\mathbb{N}$ with $t_{0}\leq t\leq\dim(F({\mathbf{x}}^{(t_{0})}))+t_{0}$ the following holds:
@@
    Finally, for $t\in\mathbb{N}$ with $t>\dim(F({\mathbf{x}}^{(t_{0})}))+t_{0}$, the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(t_{0})})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}}{8L}\cdot\left\lceil\frac{t-\dim(F({\mathbf{x}}^{(t_{0})}))-t_{0}}{m}\right\rceil\right)^{\frac{1}{2\theta-1}}.$ (3.36)
@@
Whenever Case 3 occurs we have $\dim(F({\mathbf{x}}^{(t+1)}))\leq\dim(F({\mathbf{x}}^{(t)}))-1$. On the other hand, whenever Case 1 or Case 2 occurs, the standard form description of $\mathcal{C}$ implies that $\dim(F({\mathbf{x}}^{(t+1)}))\leq\dim(F({\mathbf{x}}^{(t)}))+m$ because the reduction in the number of binding inequalities from $F({\mathbf{x}}^{(t)})$ to $F({\mathbf{x}}^{(t+1)})$ is at most $m$. Since $\dim(F({\mathbf{x}}^{(t)}))\geq 0$ for all $t\in\mathbb{N}$, it follows that for any consecutive sequence of iterations $\{t_{0},t_{0}+1,\dots,t\}$ with $t>|\dim(F({\mathbf{x}}^{(t_{0})}))|+t_{0}$ Case 1 or Case 2 must occur at least $\lceil\frac{t-\dim(F(x^{(t_{0})}))-t_{0}}{m}\rceil$ times. The rest of the proof is analogous to the proof of Theorem 3.12. ∎
@@ 3.3.2 In-face pairwise Frank-Wolfe for simplex-like polytopes
We present a FW variant that is similar to the decomposition-invariant pairwise Frank-Wolfe algorithm of Garber and Meshi [9] in Algorithm 4. Like Algorithm 3, Algorithm 4 relies on the in-face linear minimization oracle (3.21) and the maximum step-size oracle (3.22). Unlike Algorithm 3, Algorithm 4 uses the pairwise direction:
@@
Suppose that the lower bound holds for $t\in\mathbb{N}$. In case $\gamma_{t+1}\leq\eta_{t}$, Claim 3 follows from (3.28):
@@
In case $\gamma_{t+1}\geq\eta_{t}$, Claim 3 follows from the induction hypothesis and the the nonincreasing suboptimality gap:
@@
We conclude this section with the following result analogous to Proposition 3.4, Proposition 3.10, and Proposition 3.15 but specialized to the Euclidean norm. We rely on the following notation and terminology from [9]. Suppose $\mathcal{C}\subseteq\mathbb{R}^{n}$ be a simplex-like polytope and $\emptyset\neq X\subseteq\mathcal{C}$. Let $\text{\rm card}(X)$ denote the cardinality of $X$, that is, the number of positive entries of the most dense point in $X$.
@@
The bound in Proposition 3.19 is in the same spirit as the bound stated Lemma 2 in [9]. A difference is that the bound in Proposition 3.19 is sharper when the most dense point in $X^{*}$ has more than $n/2$ non-zero components. We defer the proof of Proposition 3.19 to Section 4.

### v4

@@ Abstract
We provide a template to derive affine-invariant convergence rates for the following popular versions of the Frank-Wolfe algorithm on polytopes: vanilla Frank-Wolfe, Frank-Wolfe with away steps, Frank-Wolfe with blended pairwise steps, and Frank-Wolfe with in-face directions. Our template shows how the convergence rates follow from two affine-invariant properties of the problem, namely, error bound and extended curvature. These properties depend solely on the polytope and objective function but not on any affine-dependent object like norms. For each one of the above algorithms, we derive rates of convergence ranging from sublinear to linear depending on the degree of the error bound.
@@ 1.1 Outline
The expression ${\mathcal{O}}(t^{1/(2\theta-1)})$ in the above template is to be interpreted as linear rate of convergence in the special case $\theta=1/2$. The degree $\theta\in[0,1/2]$ of the error bound property yields a rate of convergence ranging from the sublinear rate ${\mathcal{O}}(t^{-1})$ when $\theta=0$ to a linear rate when $\theta=1/2$. The main constant hidden in the ${\mathcal{O}}(t^{1/(2\theta-1)})$ expression is the ratio $L/\mu^{2\theta}$ where $L$ and $\mu$ are parameters associated to the curvature and error bound properties. As we detail in Section 2, the error bound property is defined in terms of a distance function $\mathfrak{d}:\mathcal{C}\times\mathcal{C}\rightarrow[0,1]$. The suitable distance function for each version of the Frank-Wolfe algorithm is as summarized in Table 1. The radial $\mathfrak{r}$, vertex $\mathfrak{v}$, and face $\mathfrak{f}$ distance functions are defined in Section 3.
@@ 3 Algorithms
3.1 Vanilla Frank-Wolfe algorithm.
@@
Functions of the form $\mathfrak{r}(\cdot,{\mathbf{x}})$ are somewhat intuitive and easy to visualize as they correspond to the gauge function of the set $\mathcal{C}-{\mathbf{x}}$. However, for our analysis functions of the form $\mathfrak{r}({\mathbf{y}},\cdot)$ are more relevant. Figure 4 displays visualizations of the function $\mathfrak{r}({\mathbf{y}},\cdot)$ on $\mathcal{C}$ for $\mathcal{C}=[0,1]\times[0,1]$ and various choices of ${\mathbf{y}}\in\mathcal{C}$. Figure 4(c) shows that when ${\mathbf{y}}$ is a vertex of $\mathcal{C}$, it holds that $\mathfrak{r}({\mathbf{y}},{\mathbf{x}})=1$ for every ${\mathbf{x}}\in\mathcal{C}$ with ${\mathbf{x}}\neq{\mathbf{y}}$. Figure 4(a) and Figure 4(b) show that a similar phenomenon occurs as ${\mathbf{y}}$ approaches a vertex of $\mathcal{C}$. Similar phenomena occur when ${\mathbf{y}}$ lies on, or approaches, the relative boundary of $\mathcal{C}$.
@@
For either of the line-search or short-step rules, the following progress bound holds when $\eta_{t}<1$
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle^{2}}{2L},$ (3.1)
and the following progress bound holds when $\eta_{t}=1$
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle}{2}.$ (3.2)
Our convergence analysis of FW relies on the progress bounds (3.1) and (3.2), and thus yields identical convergence rates for line-search and short-step.
@@
    $f({\mathbf{x}}^{(t)})-f^{*}\leq(f({\mathbf{x}}^{(0)})-f^{*})\left(1-\min\left\{\frac{\mu}{2L},\frac{1}{2}\right\}\right)^{t}.$ (3.3)
  • When $\theta\in[0,1/2)$ the following initial linear convergence rate holds for $t\in\{0,1,\dots,t_{0}\}$
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\frac{f({\mathbf{x}}^{(0)})-f^{*}}{2^{t}},$ (3.4)
    where $t_{0}\in\mathbb{N}$ is the smallest $t$ such that
    $\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}}{2L}\leq\frac{1}{2}.$
    Then for $t\in\mathbb{N}$ with $t>t_{0}$, the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(t_{0})})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}(t-t_{0})}{2L}\right)^{\frac{1}{2\theta-1}}.$ (3.5)
@@
Proof of Theorem 3.2. The $L$-curvature and the choice of step-size imply that for $t\in\mathbb{N}$ either (3.1) holds when $\eta_{t}<1$ or (3.2) holds when $\eta_{t}=1$. Thus, Lemma 3.3 implies that for each $t\in\mathbb{N}$ one of the following two bounds holds. If $\eta_{t}<1$ then
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle^{2}}{2L}\leq f({\mathbf{x}}^{(t)})-\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{2(1-\theta)}}{2L},$
and if $\eta_{t}=1$ then
$f({\mathbf{x}}^{(t+1)})\leq f({\mathbf{x}}^{(t)})-\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{x}}^{(t)}-{\mathbf{v}}^{(t)}\rangle}{2}\leq f({\mathbf{x}}^{(t)})-\frac{f({\mathbf{x}}^{(t)})-f^{*}}{2}.$
Hence for all $t\in\mathbb{N}$,
$f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{\mu^{2\theta}}{2L}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta},\frac{1}{2}\right\}\right).$ (3.8)
Thus (3.3) readily follows when $\theta=1/2$. On the other hand, when $\theta\in[0,1/2)$, Inequality (3.8) yields (3.4) for $t=0,1,\dots,t_{0}$. Finally for $t\in\mathbb{N}$ with $t>t_{0}$, Inequality (3.5) follows from (3.8) by applying Lemma 2.6 with $p=1-2\theta$, $\sigma_{t}=\frac{\mu^{2\theta}}{2L}$, and $\beta_{t}=f({\mathbf{x}}^{(t)})-f^{*}$. ∎
@@ 3.2 Away-step Frank-Wolfe algorithm and blended pairwise Frank-Wolfe algorithm
    For $t\in\mathbb{N}$ with $t_{0}<t\leq t_{1}:=|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$ the following holds:
@@
    Finally, for $t\in\mathbb{N}$ with $t>t_{1}$, the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(t_{1})})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}}{8L}\cdot\left\lceil\frac{t-t_{1}}{2}\right\rceil\right)^{\frac{1}{2\theta-1}}.$ (3.17)
@@
Proof of Theorem 3.6. The choice of ${\mathbf{d}}^{(t)}$ implies that for $t\in\mathbb{N}$
$\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{d}}^{(t)}\rangle\leq\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}^{(t)}-{\mathbf{a}}^{(t)}\rangle}{2}<0$ (3.20)
@@
    $f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}}{8L},\frac{1}{2}\right\}\right).$ (3.22)
@@
$f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{\mu}{8L},\frac{1}{2}\right\}\right).$
@@
Thus (3.16) holds for $t_{0}<t\leq t_{1}=|S({\boldsymbol{\lambda}}^{(t_{0})})|+t_{0}-1$ and Lemma 2.6 implies (3.17) for $t>t_{1}$. ∎
@@ 3.3 In-face Frank-Wolfe algorithm
    And for $t\in\mathbb{N}$ with $t>t_{0}$, the following sublinear convergence rate holds:
@@
Proof of Theorem 3.12. This proof is similar that of Theorem 3.6 and thus we are deliberately succinct. The choice of ${\mathbf{d}}^{(t)}$ implies that for all $t\in\mathbb{N}$
$\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{d}}^{(t)}\rangle\leq\frac{\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}^{(t)}-{\mathbf{a}}^{(t)}\rangle}{2}<0$ (3.32)
@@
    $f({\mathbf{x}}^{(t+1)})-f^{*}\leq(f({\mathbf{x}}^{(t)})-f^{*})\left(1-\min\left\{\frac{\mu^{2\theta}(f({\mathbf{x}}^{(t)})-f^{*})^{1-2\theta}}{8L},\frac{1}{2}\right\}\right).$
@@ 3.3.1 In-face FW for polytopes in standard form
    For $t\in\mathbb{N}$ with $t_{0}<t\leq t_{1}:=\dim(F({\mathbf{x}}^{(t_{0})}))+t_{0}$ the following holds:
@@
    Finally, for $t\in\mathbb{N}$ with $t>t_{1}$, the following sublinear convergence rate holds:
    $f({\mathbf{x}}^{(t)})-f^{*}\leq\left((f({\mathbf{x}}^{(t_{1})})-f^{*})^{(2\theta-1)}+\frac{(1-2\theta)\mu^{2\theta}}{8L}\cdot\left\lceil\frac{t-t_{1}}{m}\right\rceil\right)^{\frac{1}{2\theta-1}}.$ (3.39)
@@
Whenever Case 3 occurs we have $\dim(F({\mathbf{x}}^{(t+1)}))\leq\dim(F({\mathbf{x}}^{(t)}))-1$. On the other hand, whenever Case 1 or Case 2 occurs, the standard form description of $\mathcal{C}$ implies that $\dim(F({\mathbf{x}}^{(t+1)}))\leq\dim(F({\mathbf{x}}^{(t)}))+m$ because the reduction in the number of binding inequalities from $F({\mathbf{x}}^{(t)})$ to $F({\mathbf{x}}^{(t+1)})$ is at most $m$. Since $\dim(F({\mathbf{x}}^{(t)}))\geq 0$ for all $t\in\mathbb{N}$, it follows that for any consecutive sequence of iterations $\{t_{0},t_{0}+1,\dots,t\}$ with $t>t_{1}=\dim(F({\mathbf{x}}^{(t_{0})}))+t_{0}$ Case 1 or Case 2 must occur at least $\lceil\frac{t-t_{1}}{m}\rceil$ times. The rest of the proof is analogous to the proof of Theorem 3.12.∎
@@ 3.3.2 In-face pairwise Frank-Wolfe for simplex-like polytopes
We present a FW variant that is similar to the decomposition-invariant pairwise Frank-Wolfe algorithm of Garber and Meshi [9] in Algorithm 4. Like Algorithm 3, Algorithm 4 relies on the in-face linear minimization oracle (3.24). Unlike Algorithm 3, Algorithm 4 uses the pairwise direction:
@@
Suppose that the lower bound holds for $t\in\mathbb{N}_{\geq t_{0}}$. In case $\gamma_{t+1}\leq\eta_{t}$, Claim 3 follows from (3.31):
@@
In case $\gamma_{t+1}\geq\eta_{t}$, Claim 3 follows from the induction hypothesis and the nonincreasing suboptimality gap:
@@
We conclude this section with the following result analogous to Proposition 3.4, Proposition 3.10, and Proposition 3.15 but specialized to the Euclidean norm. We rely on the following notation and terminology from [9]. Suppose $\mathcal{C}\subseteq\mathbb{R}^{n}$ is a simplex-like polytope and $\emptyset\neq X\subseteq\mathcal{C}$. Let $\text{\rm card}(X)$ denote the cardinality of $X$, that is, the number of positive entries of the most dense point in $X$.
@@
The bound in Proposition 3.19 is in the same spirit as the bound stated as Lemma 2 in [9]. A difference is that the bound in Proposition 3.19 is sharper when the most dense point in $X^{*}$ has more than $n/2$ non-zero components. We defer the proof of Proposition 3.19 to Section 4.